Please help me with these problems:
Solve each system of equations by using substitution
1. x-3y=9
x+2y=-1
2. g+3h=8
1/3g+h=9
3. 2a-4b=6
-a+2b=-3
4. 2m+n=6
5m+6n=1
5. 4x-3y=-6
-x-2y=7

Answer by lagbens(14)

1.X-3Y=9 (equation 1)
X+2Y=-1 (equation 2)
let's solve for X in (eq.1):X=3Y+9
let's replace X's value in (eq.2):3Y+9+2y=-1(add up Y terms)
5Y+9=-1(subtract 9 from both side)
5Y=-10(divide both side by 5)
Y=-2
now we got Y, just replace it in one of the 2 originals equation in order to find X.
in (eq.2) we'd have X+2(-2)=-1
X-4=-1(add 4 to both side)
X=3.
final solution (3,-2)
Follow the same steps for the rest of 'em.